OpenStudy (anonymous):

State the vertical asymptote of the rational function. f(x) =(x-8)(x+4)/x^2-9

OpenStudy (noelgreco):

What's going on in the denominator of the function at a vertical asymptote?

OpenStudy (anonymous):

\[f(x)=\frac{ (x-8)(x+4) }{ x ^{2}-9 }\]

OpenStudy (anonymous):

thats the equation

OpenStudy (mertsj):

NoelGreco Medals 0 What's going on in the denominator of the function at a vertical asymptote?

OpenStudy (noelgreco):

OK, but what value does the denominator of a function have at a vertical asymptote?

OpenStudy (anonymous):

I dont know

OpenStudy (noelgreco):

The denominator is zero at a vertical asymptote. Solve x^2 - 9 =0. The solutionas are your vertical asumptotes.

OpenStudy (anonymous):

I dont get it, sorry...

OpenStudy (anonymous):

let me see if i get it

OpenStudy (mertsj):

Factor the denominator.

OpenStudy (mertsj):

What are the factors of x^2-9?

OpenStudy (anonymous):

-+ 3

OpenStudy (anonymous):

and if a letter is next to a number it means multiply

OpenStudy (mertsj):

The factors are (x-3)(x+3) Set each one equal to 0

OpenStudy (mertsj):

x-3=0---> x=3 x+3=0---x>=-3 And those are the equations of your vertical asymptotes.

OpenStudy (anonymous):

like he said that the way you do it

OpenStudy (anonymous):

dose that help out